### v1

# SupplyGraph: A Benchmark Dataset for Supply Chain Planning using Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) have gained traction across different domains such as transportation, bio-informatics, language processing, and computer vision. However, there is a noticeable absence of research on applying GNNs to supply chain networks. Supply chain networks are inherently graph-like in structure, making them prime candidates for applying GNN methodologies. This opens up a world of possibilities for optimizing, predicting, and solving even the most complex supply chain problems. A major setback in this approach lies in the absence of real-world benchmark datasets to facilitate the research and resolution of supply chain problems using GNNs. To address the issue, we present a real-world benchmark dataset for temporal tasks, obtained from one of the leading FMCG companies in Bangladesh, focusing on supply chain planning for production purposes. The dataset includes temporal data as node features to enable sales predictions, production planning, and the identification of factory issues. By utilizing this dataset, researchers can employ GNNs to address numerous supply chain problems, thereby advancing the field of supply chain analytics and planning. Source: https://github.com/CIOL-SUST/SupplyGraph
Introduction
Supply chain is a dynamic network of organizations that participate in the various processes and activities that produce value in the form of products and services for consumers via upstream and downstream linkages (Mentzer et al. 2001), entailing a continuous flow of information, goods, and money among its various stages (Higuchi and Troutt 2004). As supply chain consists of various entities interconnected in a complex network, it involves intricate inter-dependencies and complex decision-making processes. Besides, the modern supply chain produces an enormous amount of data; relationships and dependencies between all entities are not straightforward and require sophisticated models to capture (Sadeghiamirshahidi et al. 2014; Chaovalitwongse and Furman 2013).
Production planning is a crucial component of supply chain management as predicting the future demand for products or services, which helps organizations optimize their inventory levels, production schedules, and resource allocation (Aamer, Eka Yani, and Alan Priyatna 2020; Zougagh, Charkaoui, and Echchatbi 2020). Company revenue greatly relies on accurately predicting demand and planning accordingly which has led to the exploration of various deep learning and machine learning models (Alves, Silva, and Mateus 2021) (Pirhooshyaran and Snyder 2020) (Pacella and Papadia 2021) to tackle this issue. However, GNNs can model network-like structures, such as trade flows in global trade or social networks, allowing for a better understanding of complex supply chain dynamics (Kosasih and Brintrup 2022). An illustration explaining Supply chain as a graph is presented in Figure 1. Although few studies have been conducted on applying GNNs to supply chain contexts only to predict hidden link prediction to mitigate risk (Kosasih and Brintrup 2022) and uncover hidden dependency links in supply chains (Aziz et al. 2021). The absence of publicly available datasets in this domain obstructs progress and benchmarking. By furnishing a dedicated dataset, we bridge the gap between supply chain dynamics and the potential of GNNs, enabling novel models that can facilitate robust supply chain strategies for businesses. Our contributions are summarized into three folds:
  • We have introduced a new benchmark dataset for GNN problem-solving in supply chain planning, raising the bar for research and real-world applications.
  • Our dataset analysis and modeling demonstrate the superior performance of graph-based models over various SOTA approaches in diverse supply chain applications, affirming the pivotal role of GNNs in effective supply chain analysis.
Figure 1: Supply Chain Problem Formulation in Homogeneous Graph. Boxes represent various product types, with color indicating different groups. They are closely located based on product groups and production facilities. Different relational connections denote shared raw material requirements, interdependence between products, and other impacts.
Related Works
Supply Chain and Machine Learning
In the domain of production planning within supply chain management, extensive research has been dedicated to harness machine learning (Feizabadi 2022; Zhu et al. 2021; Aamer, Eka Yani, and Alan Priyatna 2020). Numerous studies aim to improve demand prediction and optimize production processes through these methodologies (Nitsche et al. 2021; Younis, Sundarakani, and Alsharairi 2022; Filali, Filali et al. 2021). For instance, deep learning techniques including Artificial Neural Networks (ANNs) (Lunardi and Lima 2021; Mrad and Boujelbene 2019; Vairagade et al. 2019), Convolutional Neural Networks (CNNs) (Husna, Amin, and Shah 2021; Tang and Ge 2022; Li 2021), and Long Short-Term Memory networks (LSTMs) (Bousqaoui, Slimani, and Achchab 2021; Lingelbach et al. 2021). Researchers have increasingly leveraged these methods to enhance the efficiency and effectiveness of production planning and supply chain management.
Table 1: Dataset Information
Node Edge Classes Edge Count
Total Products (Nodes) 41 Total Edge Types 62 Total Unique Edges 684
Number of Groups 5 Class (Group) 5 Count (Group) 188
Number of Sub-groups 19 Class (Sub-group) 19 Count (Sub-group) 52
Number of Plants 25 Class (Plant) 25 Count (Plant) 1647
Number of Storage 13 Class (Storage) 13 Count (Storage) 3046
Figure 2: Homogeneous Graphs from the dataset. (a) Nodes are sub-group products, and plants are edges. (b) Nodes plant-products and storage locations are edges. (c) Nodes are sub-group products and storage locations are edges.
Supply Chain and Graph Neural Networks
Supply Chain machine learning with Graph Neural Networks holds significant promise by enabling the modeling of complex supply chain structures, optimizing logistics, and enhancing decision-making through data-driven insights. Recent advancements in this field leverage GNNs to improve demand forecasting, and enhance supply chain resilience, paving the way for more efficient and adaptive supply chain operations. Also, the utilization of graph representation learning methodologies has led to breakthroughs in revealing hidden dependency links within supply chain networks, further elevating the performance of link prediction tasks (Aziz et al. 2021; Kosasih and Brintrup 2022) inspired by previous GNN link prediction works (Zhang and Chen 2018, 2017; Cai et al. 2020; Cai and Ji 2020). This research highlights machine learning’s effectiveness in refining demand forecasting and production planning, with GNNs showing potential in addressing supply chain challenges. However, growth in GNN-based supply chains research is very slow. It shows the significance of specialized datasets and cutting-edge methodologies for fully utilizing GNNs’ advantages in addressing the complex demand prediction and production planning issues that arise in the supply chain domain.
Data Collection
We obtained data from the central database system of one of the biggest and most prominent FMCG (Fast Moving Consumer Goods) companies in Bangladesh. We reorganize it for temporal graph utilization, extracting nodes and features. We are unable to publish the actual product names, product codes, and the name of the company, as doing so could potentially compromise the competitive standing of the concerned company. Table 1 summarizes the dataset information.
Quality Control. Each individual node, edge, and node feature undergoes a thorough manual examination and validation process. This involves scrutinizing the data for any irregularities, such as anomalies or missing information. Additionally, even zero values are meticulously reviewed to ensure their accuracy and legitimacy.
Temporal & Location Coverage. The occurrences in this database cover the period from January 1, 2023, to August 9, 2023, and include four features: production, sales orders, delivery to distributors, and factory issues in two metrics: unit and weight produced in metric tons. The company operates across the entirety of Bangladesh.
Feature Description
Within SupplyGraph, nodes pertains to distinct products, while edges represent various connections linking these products: same product group or sub-group, same plant or storage location. Figure 2 presents some examples of homogeneous graphs.
In the temporal data, node features include production, sales orders, delivery to distributors, and factory issues.
Production, which quantifies product output considering sales orders, customer demand, vehicle fill rate, and delivery urgency. This quantity is typically measured in units or Metric Tons. Sales Order signifies distributor-requested quantities, pending approval from the accounts department. It reflects overall product demand. Delivery to Distributors denotes dispatched products aligning with orders, impacting company revenue significantly. Factory Issue covers total products shipped from manufacturing facilities, with some going to distributors and the rest to storage warehouses. We have all the temporal data in two modes: number of units produced (example: 500 units) and total weight example: 10 metric tons) produced.
Data Statistics
Product Group and Sub-groups
Figure 3: Product group count.
Figure 4: Product subgroup count.
Figure 3 shows that product group “S” has the highest number of products reflecting its highest variety and product group “E” has the lowest number of products. Notably, a majority of the product groups exhibit a range of around 8 to 10 products each. Figure 4 reveals the diversity within product sub-groups. As previously noted, the "s" product group stands out for its remarkable variety, and this characteristic is consistently maintained within its corresponding sub-group. Among our product subgroups, "SOS" emerges as the one encompassing the largest array of product categories. Interestingly, despite the elevated diversity exhibited by the product group, the corresponding sub-group maintains a relatively stable count.
Analyzing Temporal Trends.
Figure 5, A particularly attractive pattern emerges: the production curve exhibits distinct and sharp fluctuations. Notably, the production unit of the company occasionally closes, which can be attributed to factors such as scheduled vacations, adjustments in the product’s MRP rate, demand variations, and optimizations in transportation policies. In contrast, the other aspects of the graph, encompassing sales orders, delivery to distributors, and factory issuance, remain notably stable and consistent. Their minimal deviations stand in contrast to the variabilities characterizing the production curve, providing an insight into the dynamics of the supply chain activities.
Figure 5: All product’s production (MT) of POV sub-group.
In, Figure 6, we can notice a very interesting pattern. The product is manufactured in large batches, with orders being accumulated over several days before production begins. The processes of delivering to distributors, handling factory issues, and fulfilling sales orders appear to be running smoothly and consistently. The company has opted to operate the production plant intermittently and maintain product inventory. When you observe straight lines in the production graph, it indicates that the plant is not actively engaged in production during those periods.
Figure 6: All 04 temporal values of product SOS008L02P.
Analyzing Temporal Correlations.
The correlation plot portrayed in Figure 7 serves to demonstrate the ways in which diverse subgroups of product "A" are interconnected in a multitude of facets. It is worth noting that the relationship between delivery to distributors and sales orders is particularly robust, whereas the production unit exhibits a distinct and prominent correlation pattern. Furthermore, a noteworthy observation pertains to the correlation between the production quantities of two subgroups, specifically "ATWWPOO2K12P" and "ATWWPOO1K24P". Additionally, it is worth highlighting that factory issues and sales orders similarly display correlation trends, albeit with some discernible variations.
Figure 7: Temporal correlation of some products from Product Group A.
Upon examining the correlation plot of the "S" subgroup in Figure 8, a distinct pattern becomes apparent: there is a significant and consistent correlation between the "SOSOO1L12P" and "SOS500M24P" subgroups across all variables. Notably, an intriguing observation arises in the context of sales orders, where a majority of the products within the "S" subgroup demonstrate a relatively higher correlation compared to delivery to distributors, factory issuances, and production.
Figure 8: Temporal correlation of some products from Product Group A.
The comprehensive analysis of correlations and temporal activities strongly indicates a substantial opportunity for the application of graph-based machine learning methodologies in the realm of supply chains. Leveraging the inherent structures and relationships within supply chain networks, graph-based machine learning holds great promise in enhancing decision-making processes, optimizing resource allocation, and fostering adaptability in the dynamic landscape of supply chain operations.
Dataset Potential and Practical Applications
Our dataset’s structural design models products as nodes, while the interrelationships, such as shared product groups, subgroups, production facilities, and storage locations, are represented as edge connections. This dataset equips us to apply GNNs to complex supply chain predicaments, such as:
Supply chain planning involves the utilization of historical data pertaining to production and demand as node attributes, with the integration of time-related features to capture seasonality and trends. To enhance demand forecasting accuracy, related entities’ influence can be harnessed through neighborhood nodes. The hierarchical structure of nodes, including product groups, subgroups, plants, and storage locations, can be used for hierarchy-aware forecasting. In data-scarce scenarios, improved forecasting can be achieved by applying transfer learning across nodes of the same classification.
Product Classification involves classifying products by product groups, sub-groups, facilities; and also by economic profitability and production similarities for economic decision-making.
Product Relation Classification involves classifying or predicting product relations in the supply chain graph. Product Relation detection involves detecting a missing edge or relation in the supply chain graph by a binary prediction objective.
Supply chain optimization involves modeling goods flow between nodes, considering demand, lead times, and reorder points. Optimal routes and quantities can be recommended using edges as transportation links. Nodes representing plants can also aid in suggesting production adjustments based on demand projections, constraints, and capacity.
Anomaly detection is achieved by contrasting predicted demand with actual data to identify disruptions, stockouts, or unusual demand spikes. Deviations in demand from aggregated neighborhood demand can signal inconsistencies within neighborhoods, which can be detected as anomalies.
Event classification involves the training of GNNs to classify events based on edges and nodes, such as production capacity changes, new product launches, and disruptions.
Fluctuation Detection entails training GNNs to identify and classify fluctuations in the network based on impacts of global price hikes, supply chain issues, and disruptions. Fluctuations in supply chains refer to disruptions that can occur due to various factors and can significantly impact the operations of businesses. These disruptions can amplify negative shocks, affecting not just the firm experiencing the failure, but also its suppliers and customers, and even firms in other parts of the production network (Acemoglu and Tahbaz-Salehi 2020). Using temporal data, the model can recognize patterns in these fluctuations over time and determine if any demand and supply fluctuation can occur in near future for better planning.
Combinatorial (Cappart et al. 2022) and Constrained (Kotary et al. 2021) Optimization involves utilizing GNNs to optimize complex decision-making within the supply chain, considering various factors and constraints. GNNs can effectively model the intricate relationships between nodes and edges, enabling efficient allocation of resources, route planning, and inventory management, ultimately leading to improved performance.
Heterogeneous Graphs. All these tasks can be executed in both heterogeneous and homogeneous forms. In the context of the heterogeneous approach, the incorporation of nodes representing distinct entities, such as products, storage locations, and production facilities, interconnected by edges that signify their intricate relationships, becomes a pivotal strategy. This approach facilitates the construction of a heterogeneous graph model, which effectively encapsulates the intricate web of interactions inherent in supply chain networks. Embracing these heterogeneous models affords us the capacity to encompass the varied characteristics of different supply chain components and their complex interdependencies. This approach proves instrumental in addressing the multifaceted challenges that characterize supply chain management. The ability to comprehensively represent and analyze these heterogeneous relationships equips decision-makers with a deeper understanding of the dynamics within the supply chain, empowering them to devise optimized strategies and adapt to the dynamic nature of supply chain operations. In doing so, overall supply chain performance and resilience can be significantly enhanced.
Hypergraphs. Efficient task execution can also be achieved through hypergraph models. Nodes, representing entities like products, storage locations, and production facilities, intricately interconnect through hyperedges, capturing the intricate relationships within supply chain networks. This strategic use of hyperedges facilitates the construction of a tailored hypergraph model, adept at encapsulating the complex web of interactions inherent in supply chain networks. The adoption of hypergraph models can provide a dynamic framework to address the diverse characteristics and intricate interdependencies of supply chain components, serving as a pivotal strategy for managing the multifaceted challenges inherent in supply chain management and fostering a deeper understanding of its dynamics, which can contribute to the enhancement of overall supply chain performance and resilience.
Experiments
Here we model our data with several benchmark graph-based models using PyTorch Geometric and PyTorch Geometric Temporal library (Rozemberczki et al. 2021). We use plant connections as edges in the graph to get predictions for all products together with 8:2 train-test split of the temporal dataset for forecasting, planning, and anomaly detection. In product classification, product relation detection, and classification, demands are used as node features. Product classification encompasses the significant task of categorizing products, while edge classification, or detection, involves discerning storage relations. Our models underwent training for 200 epochs, with the possibility of early stopping if they converged ahead of time. Different metrics were employed depending on the specific task. It’s important to note that we included non-neural network-based methods in our comparative analysis.
Experimentation and Implementation Details
Supply Chain Planning (Demand) and Forecasting. In this particular model, we adopt a temporal approach, where the node features exhibit temporal and dynamic characteristics, while nodes and edges retain static attributes to output demand (Sales Order). The edges represent the similarity between different plants or production facilities, and the products are represented as nodes. In the case of traditional models, we independently model each product and subsequently aggregate the results. Conversely, with temporal Graph Neural Networks (GNNs), all predictions are generated simultaneously. It’s worth noting that we have maintained the use of default hyperparameters, as mentioned previously. MSE loss is used.
Among the models, ARMA (Reinsel 1993) is a traditional statistical model, GRU (Chung et al. 2014) and LSTM (Hochreiter and Schmidhuber 1997) are deep learning based models, DCRNN (Li et al. 2018), TGCN (Zhao et al. 2018), GConvGRU (Seo et al. 2016) are GNN-based models. Our approach for production forecasting model is same to demand prediction model, with the distinction being that the primary output objective was production rather than demand.
Product Classification. In the context of product classification, we harnessed the temporal features as static node attributes to forecast the product group, with the edges denoting the similarities between production facilities. The loss function was cross-entropy loss. Our Artificial Neural Network (ANN) model featured two linear layers, the K-Nearest Neighbors (KNN) algorithm incorporated a setting of 5 closest neighbors, and the Graph Attention Network (GAT) employed 4 attention heads with 3 layers in each Graph Neural Network (GNN). Amog the models, Logistic Regression is a statistical model, XGB Classifier (Chen and Guestrin 2016) is also a traditional bossting based model, KNN (Mucherino, Papajorgji, and Pardalos 2009) is a statistical clustering based model, ANN is a deep learning based model, GCN (Kipf and Welling 2017) and GAT (Veličković et al. 2018) are GNN-based models.
Anomaly Detection. For anomaly detection, we employed sales orders as dynamic temporal node characteristics, applying a model with Mean Squared Error (MSE) loss to identify anomalies within time series data. Increased data fluctuations across the graph during a specific time frame are anomaly as an anomaly and warrant communication with the company. It is a supervised approach; following GDN (Deng and Hooi 2021). Among the models, PCA is a statistical model; ANN, AE (Aggarwal 2015) and LSTM-VAE (Park, Hoshi, and Kemp 2017) are deep learning based models; GANF (Dai and Chen 2022) and GDN (Deng and Hooi 2021) are GNN-based temporal anomaly detection models.
Product Relation Detection and Classification. Product relation detection entailed a binary classification task focused on discerning the presence or absence of edges, while product relation classification involved a multi-class classification task. The loss function utilized in this context was the cross-entropy loss. As before, the ANN model featured two linear layers, the K-Nearest Neighbors (KNN) algorithm incorporated a setting of 5 closest neighbors, and the Graph Attention Network (GAT) employed 4 attention heads with 3 layers in each Graph Neural Network (GNN).
Model RMSE MAE
ARMA 37.9875 486.8858
GRU 33.2277 359.8285
LSTM 30.7782 304.6777
DCRNN 27.9811 253.1135
TGCN 28.0145 253.9384
GConvGRU 28.1814 272.6487
Table 2: Supply Chain Planning (Demand)
Model RMSE MAE
ARMA 37.4542 483.5387
GRU 32.3532 354.2312
LSTM 30.2142 300.2942
DCRNN 27.9434 252.6459
TGCN 27.9244 253.2390
GConvGRU 28.0414 270.5629
Table 3: Production Forecasting
Model Accuracy Precision
Logistic Regression 66.67% 76.67%
XGB Classifier 55.56% 41.48%
KNN 44.44% 74.07%
ANN 66.67% 77.78%
GCN 61.12% 72.18%
GAT 65.92% 75.32%
Table 4: Classification
Model Precision Recall
PCA 31.33% 27.34%
ANN 55.24% 47.75%
AE 75.78% 71.78%
LSTM-VAE 91.87% 86.58%
GANF 85.45% 76.33%
GDN 94.15% 87.76%
Table 5: Anomaly Detection
Model Accuracy Precision
Logistic Reg. 68.63% 75.15%
XGB Classifier 71.23% 81.12%
KNN 74.63% 78.26%
ANN 78.74% 81.62%
GCN 89.32% 88.28%
GAT 91.45% 91.64%
Table 6: Product Relation Detection
Model Accuracy Precision
Logistic Reg. 62.73% 66.78%
XGB Classifier 71.78% 71.86%
KNN 54.72% 74.75%
ANN 76.58% 77.72%
GCN 91.36% 82.27%
GAT 89.43% 81.22%
Table 7: Product Relation Classification
Results
Table 2-7 reveals that graph-based models consistently achieve higher scores compared to other models in the experiments. The experiments were performed multiple times and an average is taken.
Supply Chain Planning (Demand). In demand prediction (see Table 2), we observe a range of models, including traditional time series methods like ARMA and more advanced recurrent neural networks such as GRU and LSTM. Notably, the newer graph-based Graph Convolutional GRU (GConvGRU) achieves competitive performance. It’s evident that the recurrent models, particularly DCRNN and TGCN, outperform ARMA, demonstrating the benefits of leveraging sequential information with GNNs. GNN-based models like GConvGRU also showcase their potential in modeling supply chain demand patterns.
Production Forecasting Similar to demand prediction (see Table 3), production forecasting shows a similar trend with respect to model performance. The recurrent models, DCRNN and TGCN, continue to outperform ARMA, emphasizing the superiority of GNN-based methods for this task. LSTM stands out as the best-performing individual model, closely followed by DCRNN and TGCN, signifying their utility in forecasting production levels using GNNs.
Product Classification In product classification demostrated in Table 4, traditional logistic regression achieves a respectable accuracy of 66.67%, but GNN-based models like GCN and GAT exhibit competitive performance with an accuracy of 61.12% and 65.92%, respectively. The hybrid models, such as GConvGRU and TGCN, offer promise in classifying products. The performance of these models showcases their adaptability for supply chain product classification tasks.
Anomaly Detection Anomaly detection involves identifying deviations from expected patterns in time series. In this task, Table 5 shows that LSTM-VAE and GDN lead the way in terms of precision and recall, underlining their effectiveness in pinpointing anomalies. The GANF model also demonstrates strong performance, emphasizing the potential of GNN-based approaches for anomaly detection.
Product Relation Detection The task of detecting relationships between products is approached effectively by various models. Table 6 shows that, GCN and GAT outperform the traditional approaches, achieving an accuracy of 89.32% and 91.45%, respectively. These results underscore the capacity of GNNs to model intricate product relationships in the supply chain.
Product Relation Classification In product relation classification demostrated in Table 7, showcases the efficiency of GNNs, with GCN and GAT achieving remarkable accuracy levels of 91.36% and 89.43%, respectively. They significantly outperform traditional methods and even other machine learning algorithms, demonstrating their prowess in classifying complex product relationships.
Discussion
We firmly believe that this dataset can serve as a foundational cornerstone for supply chain analysis and problem-solving using Graph Neural Networks (GNNs). In the current landscape, a significant limitation has been the lack of proper datasets tailored to the complexities of supply chain management. The importance of GNN-based supply chain management cannot be overstated, as it allows for the modeling of intricate network relationships, leading to more robust decision-making and optimization. Furthermore, the development of a benchmark graph dataset specifically designed for supply chain applications, covering various tasks and challenges, is crucial for advancing the field and promoting standardized evaluation metrics and methodologies in this domain.
We acknowledge that the current temporal data is limited by the timeframe. Expanding the temporal scope of available data would undoubtedly provide valuable benefits by offering a more extensive vantage point on longer-term supply chain dynamics. Despite its temporal brevity, the results show that it remains a valuable asset for the examination and resolution of diverse supply chain challenges. Its relatively concise time frame proves suitable for the assessment of immediate responses to events and decisions, providing valuable insights into the short-term intricacies of supply chain dynamics.
The utility of the dataset hinges on the choice of the target model or the problem to be addressed, and it is subject to various perspectives. By considering different pieces of information as nodes or edges, the dataset’s potential applications and implications can be substantial. Ultimately, it is at the researcher to determine how to leverage it effectively.
Conclusion
Supply chain networks present a potent framework for optimization and problem-solving for GNNs. However, the lack of real-world benchmark datasets impedes progress in this field. To address this issue, our work aims to provide a real-world benchmark graph-based dataset from a prominent FMCG firm in Bangladesh, with a focus on production-oriented supply chain planning. This dataset includes interconnected product nodes, as well as temporal features for sales prediction, production planning, and issue identification. Experimental results establish the importance of GNN-based methods in supply chains. Researchers can utilize this dataset to leverage graph neural networks in resolving multifaceted supply chain issues, thus catalyzing advancements in supply chain analytics and planning.
Acknowledgement
We thank the reviewers for their thoughtful comments and feedback. We appreciate all the support from the Computational Intelligence and Operations Lab, SUST.